OpenStudy (anonymous):

The first term of an arithmetic sequence is 2 and the fourth term is 11. Find the sum of the first 50 terms.

OpenStudy (anonymous):

3775

OpenStudy (anonymous):

how did u find that number?

OpenStudy (anonymous):

The nth number in an arithmetic sequence = a+(n-1)d where a is the first term and d is the difference The sum of the first n numbers in an A.P. is n/2(a+nth number)

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

The sum of 50 terms = 50/2 [2a + (n-1)d] 11 = 2 + 3d 9 = 3d 9/3 = d 3 = d sum of 50 terms = 50/2 [2a + (n-1)d] = 25[2*2 + 49*3] = 3775